# Plan-R1: Safe and Feasible Trajectory Planning as Language Modeling

### v1

@@ Abstract
Safe and feasible trajectory planning is essential for real-world autonomous driving systems. However, existing learning-based planning methods often rely on expert demonstrations, which not only lack explicit safety awareness but also risk inheriting unsafe behaviors such as speeding from suboptimal human driving data. Inspired by the success of large language models, we propose Plan-R1, a novel two-stage trajectory planning framework that formulates trajectory planning as a sequential prediction task, guided by explicit planning principles such as safety, comfort, and traffic rule compliance. In the first stage, we train an autoregressive trajectory predictor via next motion token prediction on expert data. In the second stage, we design rule-based rewards (e.g., collision avoidance, speed limits) and fine-tune the model using Group Relative Policy Optimization (GRPO), a reinforcement learning strategy, to align its predictions with these planning principles. Experiments on the nuPlan benchmark demonstrate that our Plan-R1 significantly improves planning safety and feasibility, achieving state-of-the-art performance.
@@ 3.2 Autoregressive pre-training
Tokenization. To facilitate sequence modeling, continuous trajectories are discretized into motion tokens. Temporally, trajectories are segmented at fixed intervals. Spatially, the K-disk clustering algorithm [23] is applied to the resulting motion segments based on their average corner distance, producing a motion token vocabulary. Each token represents a prototypical displacement and heading change over a fixed time step. For each agent, the entire trajectory is thus transformed into a sequence of such motion tokens spanning the prediction horizon.
Model architecture. Our model employs a transformer decoder with factorized attention [19, 20] to encode multi-agent spatiotemporal interactions, as shown in Figure 1 (a). It is designed to capture the temporal dynamics of each agent’s trajectory and its interactions with the map and nearby agents.
@@
Training objective. The model is trained with a next motion token prediction objective. Given a tokenized trajectory sequence, the model is trained to minimize the negative log-likelihood of the next motion token at each step:
@@ 3.3 Reinforcement learning fine-tuning
Dual-Model rollout. To enable stable multi-agent rollout and principle-aligned fine-tuning, we adopt a dual-model design. Specifically, we freeze the pretrained autoregressive model for predicting the future motions of the surrounding agents, and create a trainable replica to represent the ego policy $\pi_{e}$. During each rollout, the two models jointly generate trajectories step by step. The frozen model, serving as a reactive world model, predicts the future motions of surrounding agents based on the history of all agents, including the ego. Although fixed during training, the world model remains responsive to ego actions by updating its predictions based on the ego’s historical motion tokens. This interaction ensures that the surrounding agents’ behaviors are aligned with the ego vehicle’s actions, ensuring realistic, interaction-aware simulation in multi-agent settings. Meanwhile, the trainable ego model explores alternative actions and receives feedback through the reward function. This separation enables ego-centric policy updates without destabilizing surrounding agent dynamics, while also ensuring that all agents exhibit natural reactive behaviors, as observed in real-world interactions.
Rule-based rewards. To align the ego policy with planning principles, we design a set of interpretable, rule-based reward functions. Rewards are computed based on the final predicted trajectories and are designed to reflect key planning properties such as safety, feasibility, and comfort. Following [2, 9], we compute the total reward as the product of two components: a set of multiplicative safety indicators (e.g., driving area compliance) and a weighted sum of soft cost terms (e.g., comfort):
@@
where $\mathbf{1}_{k}\in\{0,1\}$ denotes whether safety constraint $k$ is satisfied, and $r_{j}(Y)$ is the score of cost term $j$ with weight $w_{j}$. This formulation ensures that violations of critical safety conditions will nullify the total reward, while soft planning objectives are optimized when safety constraints are satisfied. Details of the reward components and implementation are provided in the Appendix.
Optimization with GRPO. To fine-tune the ego policy toward planning principles, we adopt GRPO as our reinforcement learning algorithm. Compared to PPO, GRPO eliminates the need for an explicit value function by computing relative advantages within sampled trajectory groups. This reduces implementation complexity, avoids the instability associated with inaccurate value estimates, and improves sampling efficiency, making it well-suited for trajectory-level planning.
@@ 4.1 Experimental setup
Datasets. We evaluate our method on the nuPlan [2] benchmark, a large-scale platform designed for trajectory planning in autonomous driving. The dataset contains over 1,300 hours of expert driving logs collected across four cities, covering diverse real-world scenarios, such as merges, intersections, roundabouts, lane changes, and interactions with pedestrians. The benchmark supports simulation-based evaluation, where scenarios are replayed for 15 seconds at 10 Hz. The ego vehicle executes planned trajectories via a bicycle model and an LQR controller, while other traffic agents follow either replayed or Intelligent Driver Model (IDM) [33] policies. Following PLUTO [4], we sample 1M training instances across all scenario types for autoregressive pretraining. To reduce the computational cost of closed-loop rollouts during reinforcement learning, we construct a smaller dataset of 100,000 scenarios for fine-tuning. Evaluations are conducted on the standard Val14 [6], Test14-random, and Test14-hard [5] splits under both non-reactive and reactive simulation settings.
Metrics. We evaluate planning performance in closed-loop simulation using the Non-Reactive Closed-Loop Score (NR-CLS) and Reactive Closed-Loop Score (R-CLS). These metrics assess the planner’s ability to generate safe and feasible trajectories over 15-second rollouts, considering collisions, off-road driving, speed limit violations, and the progress of the expert route. In the NR-CLS setting, surrounding agents follow replayed logged trajectories, while in R-CLS, they react using the IDM [33] planner. Each score ranges from 0 to 100, where a higher score indicates better performance.
@@ 4.2 Comparison with SOTAs
Quantitative results. We compare our method, Plan-R1, against a wide range of existing planners on the nuPlan benchmark, as shown in Table 1. Across all three benchmarks, Plan-R1 achieves performance comparable to the strongest prior method, Diffusion Planner [39], under the non-reactive setting. This indicates that our reinforcement learning fine-tuning preserves the model’s ability to align with expert-like behavior. Under the more challenging reactive setting, where surrounding agents respond dynamically to the ego vehicle, Plan-R1 substantially outperforms all baselines. It achieves state-of-the-art reactive scores of 84.97 on Val14, 73.18 on Test14-hard, and 94.51 on Test14-random, surpassing the next-best method, Diffusion Planner, by +2.17, +3.96, and +5.96 points, respectively. Compared to other learning-based planners such as PlanTF [5], Plan-R1 demonstrates significantly stronger generalization to interactive scenarios. We attribute this improvement to the exploration capability introduced by reinforcement learning, which enables the planner to generate safer and more feasible trajectories beyond what is directly covered in the expert demonstrations. Following Diffusion Planner [39], we apply the existing refinement module [30] as a post-processing step to Plan-R1, ensuring a fair comparison with prior methods that adopt similar pipelines. With post-processing, Plan-R1 continues to achieve state-of-the-art performance, reaching a reactive score of 94.51 on Test14-random, outperforming the second-best method by 2.76 points. This result highlights the effectiveness of our approach in producing high-quality trajectories.
@@
Qualitative Results. To further illustrate the practical benefits of our approach, we conduct a case study comparing closed-loop non-reactive rollouts across different planners on a representative scenario from the nuPlan dataset. In this scenario, the expert trajectory exhibits a speeding behavior that violates the local speed limit. We evaluate three learning-based planners—PLUTO [4], Diffusion Planner [39], and Plan-R1—none of which incorporate rule-based post-processing in this analysis. As shown in Figure 2, both PLUTO and Diffusion Planner, which are trained solely via supervised learning on expert demonstrations, replicate the speeding behavior in their closed-loop rollouts. In contrast, Plan-R1 successfully avoids speeding and maintains a compliant velocity profile throughout the rollout. This demonstrates our model’s ability to identify and correct undesirable behaviors inherited from expert data, a capability enabled by reinforcement learning fine-tuning with rule-based reward supervision. These results highlight the importance of explicit objective alignment in ensuring rule-compliant and safe planning behavior.
@@ 4.3 Ablation studies
The importance of the fine-tuning stage. To evaluate the effectiveness of the fine-tuning stage in our Plan-R1 framework, we compare Plan-R1 with a baseline model that undergoes only pretraining, as shown in Table 2. Compared to the baseline without fine-tuning, Plan-R1 shows substantial improvements across key metrics in the reactive closed-loop simulation on Test14-random. Specifically, it achieves a score increase of +6.08, with notable gains in collision avoidance (+2.11), time-to-collision (TTC) (+4.21), drivable area compliance (+3.06), and speed limit compliance (+1.63). Although the comfort score fluctuated slightly (-0.16), the progress score improved significantly (+2.98), suggesting that the model enhances safety and feasibility not by relying on remaining stationary or abrupt braking, but by generating better trajectories.
@@
Effect of KL regularization. We investigate the impact of the KL divergence regularization weight $\beta$ in Equation 5 during the fine-tuning stage. As shown in Figure 5, when $\beta$ is set too small (e.g., 0.01), the updated policy deviates excessively from the pretrained reference, leading to a degradation in performance due to the loss of learned motion priors. Conversely, when $\beta$ is too large (e.g., 1.0), the policy is overly constrained to remain close to the reference model, limiting the benefits of fine-tuning and resulting in similar results. A moderate value (e.g., $\beta{=}0.1$) achieves the best balance, preserving the distributional prior while enabling meaningful adaptation toward planning objectives.
Pass@$k$ analysis. To further assess the reliability of our method under varying sampling conditions, we conduct a pass@$k$ [14] analysis on the validation set. For both the pretrained baseline and our Plan-R1, we generate $k$ trajectories per scenario and compute the proportion of cases where at least one satisfies all core safety and feasibility criteria—i.e., staying within drivable areas, avoiding collisions, respecting speed limits, and maintaining comfort. As shown in Figure 5, pass@$k$ improves with increasing $k$ for both models. However, Plan-R1 consistently outperforms the baseline, indicating it yields high-quality trajectories with greater likelihood. This suggests that fine-tuning not only improves the single-shot performance but also enhances the model’s ability to generate feasible trajectories across multiple samples, which is critical for robust planning under uncertainty.

### v2

@@ Abstract
Safe and feasible trajectory planning is essential for real-world autonomous driving systems. However, existing learning-based planning methods often rely on expert demonstrations, which not only lack explicit safety awareness but also risk inheriting unsafe behaviors such as speeding from suboptimal human driving data. Inspired by the success of large language models, we propose Plan-R1, a novel two-stage trajectory planning framework that formulates trajectory planning as a sequential prediction task, guided by explicit planning principles such as safety, comfort, and traffic rule compliance. In the first stage, we train an autoregressive trajectory predictor via next motion token prediction on expert data. In the second stage, we design rule-based rewards (e.g., collision avoidance, speed limits) and fine-tune the model using Group Relative Policy Optimization (GRPO), a reinforcement learning strategy, to align its predictions with these planning principles. Experiments on the nuPlan benchmark demonstrate that our Plan-R1 significantly improves planning safety and feasibility, achieving state-of-the-art performance. Our code will be made public soon.
@@ 3.2 Autoregressive pre-training
Tokenization.
To facilitate sequence modeling, continuous trajectories are discretized into motion tokens. Temporally, trajectories are segmented at fixed intervals. Spatially, the K-disk clustering algorithm [24] is applied to the resulting motion segments based on their average corner distance, producing a motion token vocabulary. Each token represents a prototypical displacement and heading change over a fixed time step. For each agent, the entire trajectory is thus transformed into a sequence of such motion tokens spanning the prediction horizon.
Model architecture.
Our model employs a transformer decoder with factorized attention [20, 21] to encode multi-agent spatiotemporal interactions, as shown in Figure 1 (a). It is designed to capture the temporal dynamics of each agent’s trajectory and its interactions with the map and nearby agents.
@@
Training objective.
The model is trained with a next motion token prediction objective. Given a tokenized trajectory sequence, the model is trained to minimize the negative log-likelihood of the next motion token at each step:
@@ 3.3 Reinforcement learning fine-tuning
Dual-Model rollout.
To enable stable multi-agent rollout and principle-aligned fine-tuning, we adopt a dual-model design. Specifically, we freeze the pretrained autoregressive model for predicting the future motions of the surrounding agents, and create a trainable replica to represent the ego policy $\pi_{e}$. During each rollout, the two models jointly generate trajectories step by step. The frozen model, serving as a reactive world model, predicts the future motions of surrounding agents based on the history of all agents, including the ego. Although fixed during training, the world model remains responsive to ego actions by updating its predictions based on the ego’s historical motion tokens. This interaction ensures that the surrounding agents’ behaviors are aligned with the ego vehicle’s actions, ensuring realistic, interaction-aware simulation in multi-agent settings. Meanwhile, the trainable ego model explores alternative actions and receives feedback through the reward function. This separation enables ego-centric policy updates without destabilizing surrounding agent dynamics, while also ensuring that all agents exhibit natural reactive behaviors, as observed in real-world interactions.
Rule-based rewards.
To align the ego policy with planning principles, we design a set of interpretable, rule-based reward functions. Rewards are computed based on the final predicted trajectories and are designed to reflect key planning properties such as safety, feasibility, and comfort. Following [2, 9], we compute the total reward as the product of two components: a set of multiplicative safety indicators (e.g., driving area compliance) and a weighted sum of soft cost terms (e.g., comfort):
@@
where $\mathbf{1}_{k}\in\{0,1\}$ denotes whether safety constraint $k$ is satisfied, and $r_{j}(Y)$ is the score of cost term $j$ with weight $w_{j}$. This formulation ensures that violations of critical safety conditions will nullify the total reward, while soft planning objectives are optimized when safety constraints are satisfied. Details of the reward components and implementation are provided in the Section A.2.
Optimization with GRPO.
To fine-tune the ego policy toward planning principles, we adopt GRPO as our reinforcement learning algorithm. Compared to PPO, GRPO eliminates the need for an explicit value function by computing relative advantages within sampled trajectory groups. This reduces implementation complexity, avoids the instability associated with inaccurate value estimates, and improves sampling efficiency, making it well-suited for trajectory-level planning.
@@ 4.1 Experimental setup
Datasets.
We evaluate our method on the nuPlan [2] benchmark, a large-scale platform designed for trajectory planning in autonomous driving. The dataset contains over 1,300 hours of expert driving logs collected across four cities, covering diverse real-world scenarios, such as merges, intersections, roundabouts, lane changes, and interactions with pedestrians. The benchmark supports simulation-based evaluation, where scenarios are replayed for 15 seconds at 10 Hz. The ego vehicle executes planned trajectories via a bicycle model and an LQR controller, while other traffic agents follow either replayed or Intelligent Driver Model (IDM) [34] policies. Following PLUTO [4], we sample 1M training instances across all scenario types for autoregressive pretraining. To reduce the computational cost of closed-loop rollouts during reinforcement learning, we construct a smaller dataset of 100,000 scenarios for fine-tuning. Evaluations are conducted on the standard Val14 [6], Test14-random, and Test14-hard [5] splits under both non-reactive and reactive simulation settings.
Metrics.
We evaluate planning performance in closed-loop simulation using the Non-Reactive Closed-Loop Score (NR-CLS) and Reactive Closed-Loop Score (R-CLS). These metrics assess the planner’s ability to generate safe and feasible trajectories over 15-second rollouts, considering collisions, off-road driving, speed limit violations, and the progress of the expert route. In the NR-CLS setting, surrounding agents follow replayed logged trajectories, while in R-CLS, they react using the IDM [34] planner. Each score ranges from 0 to 100, where a higher score indicates better performance.
@@ 4.2 Comparison with SOTAs
Quantitative results.
We compare our method, Plan-R1, against a wide range of existing planners on the nuPlan benchmark, as shown in Table 1. Across all three benchmarks, Plan-R1 achieves performance comparable to the strongest prior method, Diffusion Planner [40], under the non-reactive setting. This indicates that our reinforcement learning fine-tuning preserves the model’s ability to align with expert-like behavior. Under the more challenging reactive setting, where surrounding agents respond dynamically to the ego vehicle, Plan-R1 substantially outperforms all baselines. It achieves state-of-the-art reactive scores of 84.97 on Val14, 73.18 on Test14-hard, and 94.51 on Test14-random, surpassing the next-best method, Diffusion Planner, by +2.17, +3.96, and +5.96 points, respectively. Compared to other learning-based planners such as PlanTF [5], Plan-R1 demonstrates significantly stronger generalization to interactive scenarios. We attribute this improvement to the exploration capability introduced by reinforcement learning, which enables the planner to generate safer and more feasible trajectories beyond what is directly covered in the expert demonstrations. Following Diffusion Planner [40], we apply the existing refinement module [31] as a post-processing step to Plan-R1, ensuring a fair comparison with prior methods that adopt similar pipelines. With post-processing, Plan-R1 continues to achieve state-of-the-art performance, reaching a reactive score of 94.51 on Test14-random, outperforming the second-best method by 2.76 points. This result highlights the effectiveness of our approach in producing high-quality trajectories.
@@
Qualitative Results.
To further illustrate the practical benefits of our approach, we conduct a case study comparing closed-loop non-reactive rollouts across different planners on a representative scenario from the nuPlan dataset. In this scenario, the expert trajectory exhibits a speeding behavior that violates the local speed limit. We evaluate three learning-based planners—PLUTO [4], Diffusion Planner [40], and Plan-R1—none of which incorporate rule-based post-processing in this analysis. As shown in Figure 2, both PLUTO and Diffusion Planner, which are trained solely via supervised learning on expert demonstrations, replicate the speeding behavior in their closed-loop rollouts. In contrast, Plan-R1 successfully avoids speeding and maintains a compliant velocity profile throughout the rollout. This demonstrates our model’s ability to identify and correct undesirable behaviors inherited from expert data, a capability enabled by reinforcement learning fine-tuning with rule-based reward supervision. These results highlight the importance of explicit objective alignment in ensuring rule-compliant and safe planning behavior.
@@ 4.3 Ablation studies
The importance of the fine-tuning stage.
To evaluate the effectiveness of the fine-tuning stage in our Plan-R1 framework, we compare Plan-R1 with a baseline model that undergoes only pretraining, as shown in Table 2. Compared to the baseline without fine-tuning, Plan-R1 shows substantial improvements across key metrics in the reactive closed-loop simulation on Test14-random. Specifically, it achieves a score increase of +6.08, with notable gains in collision avoidance (+2.11), time-to-collision (TTC) (+4.21), drivable area compliance (+3.06), and speed limit compliance (+1.63). Although the comfort score fluctuated slightly (-0.16), the progress score improved significantly (+2.98), suggesting that the model enhances safety and feasibility not by relying on remaining stationary or abrupt braking, but by generating better trajectories.
@@
Effect of KL regularization.
We investigate the impact of the KL divergence regularization weight $\beta$ in Equation 5 during the fine-tuning stage. As shown in Figure 5, when $\beta$ is set too small (e.g., 0.01), the updated policy deviates excessively from the pretrained reference, leading to a degradation in performance due to the loss of learned motion priors. Conversely, when $\beta$ is too large (e.g., 1.0), the policy is overly constrained to remain close to the reference model, limiting the benefits of fine-tuning and resulting in similar results. A moderate value (e.g., $\beta{=}0.1$) achieves the best balance, preserving the distributional prior while enabling meaningful adaptation toward planning objectives.
Pass@$k$ analysis.
To further assess the reliability of our method under varying sampling conditions, we conduct a pass@$k$ [14] analysis on the validation set. For both the pretrained baseline and our Plan-R1, we generate $k$ trajectories per scenario and compute the proportion of cases where at least one satisfies all core safety and feasibility criteria—i.e., staying within drivable areas, avoiding collisions, respecting speed limits, and maintaining comfort. As shown in Figure 5, pass@$k$ improves with increasing $k$ for both models. However, Plan-R1 consistently outperforms the baseline, indicating it yields high-quality trajectories with greater likelihood. This suggests that fine-tuning not only improves the single-shot performance but also enhances the model’s ability to generate feasible trajectories across multiple samples, which is critical for robust planning under uncertainty.
@@ 5 Conclusion
Appendix A Technical Appendices and Supplementary Material
A.1 Additional ablation study: parameter scaling baseline
Since our Plan-R1 integrates a pretrained trajectory predictor and a fine-tuned planner, it potentially increases the overall parameter count. To assess whether the performance improvement of Plan-R1 stems purely from this increase in model capacity, we conduct an ablation study comparing it with a scaled-up baseline model. This scaled baseline increases the hidden dimension from 128 to 192, resulting in approximately doubling the number of parameters.
As shown in Table 3, the Baseline (2$\times$) shows modest improvements over the original baseline in a few metrics, such as overall score (+2.13), collision avoidance (+1.15), and progress (+1.40). However, several critical metrics, including TTC, speed limit compliance, and comfort, do not improve with increased capacity; in fact, their performance slightly degrades (TTC: -0.38, Speed: -0.24, Comfort: -1.15). We attribute this phenomenon to the imperfections in the expert data, which contain suboptimal behaviors such as speeding, uncomfortable motion patterns, and overly short TTC. As the model capacity increases, the baseline becomes more capable of faithfully imitating these undesirable patterns, leading to the reproduction—and even amplification—of unsafe or uncomfortable behaviors in simulation.
In contrast, Plan-R1 achieves consistent and substantial improvements across almost all metrics. Compared to the Baseline (2$\times$), it improves the overall score by +3.95 and achieves notable gains in TTC (+4.59), speed limit compliance (+1.87), and other safety-related metrics. These results indicate that Plan-R1 benefits not only from increased capacity, but more importantly from its structured design. In particular, the fine-tuning stage explicitly aligns generation with planning principles via rule-based rewards, mitigating undesirable behaviors inherited from expert data and improving safety awareness in trajectory generation.
Table 3: Comparison between the original baseline, a scaled-up baseline with approximately twice the parameter count (Baseline 2$\times$), and Plan-R1.
Planner Score Collisions TTC Drivable Speed Comfort Progress
Baseline 82.81 94.44 90.04 94.64 97.83 99.62 85.71
Baseline (2$\times$) 84.94 95.59 89.66 95.79 97.59 98.47 87.11
Plan-R1 88.89 96.55 94.25 97.70 99.46 99.46 88.69
A.2 Reward design
We design the total reward as a product of two components: a set of binary safety indicators and a weighted sum of soft cost terms. This formulation first ensures that no reward is given if any critical safety condition is violated, and then encourages the agent to optimize for smoother and more efficient trajectories when safety is satisfied. Below, we describe the components in detail.
Safety indicators ($\mathcal{I}_{\text{safe}}$).
The following constraints must all be satisfied to receive a non-zero reward:
  • Driving area compliance: All predicted positions must lie within the drivable area polygon.
  • No collision with dynamic agents: At each prediction step, the ego vehicle must not collide with the predicted positions of any surrounding agent at the corresponding time step.
  • No collision with static obstacles: All predicted positions must avoid collisions with static obstacles.
To ensure accurate safety assessment, all collision checks are performed using the full bounding boxes of the ego vehicle and surrounding agents, rather than using center points.
Soft cost terms ($\mathcal{I}_{\text{cost}}$).
When all safety indicators are satisfied, the agent receives a soft reward computed as a weighted sum of four normalized terms:
  • Comfort ($w=2$): This term penalizes high lateral and longitudinal accelerations, angular velocity, and angular acceleration. If any of these exceed predefined thresholds, the comfort score is set to 0; otherwise, it is 1.
  • Time-to-collision (TTC) ($w=4$): This term is set to 1 if the predicted TTC remains above a safety threshold throughout the trajectory; otherwise, it is set to 0.
  • Speed limit compliance ($w=5$): This term linearly penalizes the predicted speeding behavior by computing the ratio between the total over-speed distance and a predefined maximum over-speed threshold. The score decreases linearly from 1 to 0 as this ratio increases, following the same formulation as in the nuPlan benchmark [2].
  • Progress ($w=2$): This term measures the forward progress of the predicted trajectory along the reference route. It is normalized using the progress of the expert trajectory. The progress is computed by projecting each predicted position onto the centerline of the reference route and accumulating the forward distance along the projected path. Note that this is not an imitation loss: we do not minimize the distance between the predicted and expert trajectories, but only use the expert’s final progress as a normalization reference.
Each reward term $r_{j}(Y)$ is normalized to the range $[0,1]$, and the overall soft reward is computed as a weighted sum $\sum_{j\in\mathcal{I}_{\text{cost}}}w_{j}\cdot r_{j}(Y)$. This design encourages smooth, efficient, and rule-compliant trajectories, while ensuring that any violation of critical safety conditions suppresses the entire reward.
A.3 Implementation details
Architecture.
We use a discrete motion token vocabulary of size 1024 for each agent category (Vehicle, Pedestrian, Cyclist), where each token corresponds to a 0.5-second movement segment. The ego agent shares the same token vocabulary as other vehicles in the Vehicle category. Our model adopts a 6-layer Transformer decoder, each with 8 attention heads and a hidden dimension of 128. At the end of each decoder layer, a token prediction head is applied, implemented as a two-layer MLP. The parameters of an autoregressive decoder are approximately 5.05M.
Training.
During pretraining, we use teacher forcing to optimize next motion token prediction with cross-entropy loss. The model is trained for 32 epochs with a batch size of 64 using the AdamW optimizer [19], with a learning rate of $3\times 10^{-4}$ and a weight decay of $1\times 10^{-4}$. A dropout rate of 0.1 is applied throughout training. A cosine annealing scheduler is used to decay the learning rate to zero during pretraining. For fine-tuning, we apply GRPO with a group size of $G=4$, KL divergence regularization weight $\beta=0.1$, and a reduced learning rate of $4\times 10^{-5}$. Fine-tuning is performed for 5 epochs, using the same dropout rate of 0.1 as in pretraining. No data augmentation is applied during either training or fine-tuning. All experiments are conducted on 8 NVIDIA RTX 4090 GPUs.
Simulation.
During simulation, both the pretrained and fine-tuned models generate trajectories by selecting the top-1 token at each step.
A.4 Limitations and future work
The fine-tuning stage of our method relies on a set of rule-based rewards, whose design and weight calibration require substantial manual effort and iterative tuning. Furthermore, the fine-tuning process depends on trajectories sampled from the pretrained model, which may fail to capture rare yet critical behavior,such as emergency braking to avoid collisions, due to their limited presence in expert demonstrations. As part of future work, we plan to collect a set of safety-critical scenarios and apply supervised fine-tuning (SFT) to provide a stronger initialization prior to reinforcement learning.
A.5 Broader impacts
This work introduces a novel perspective on trajectory planning by framing it as a principle-aligned autoregressive generation task. By decoupling behavior learning from planning principle alignment, our approach facilitates the generation of safer and more feasible trajectories in complex driving scenarios. The proposed framework has the potential to advance the development of more reliable planning systems for autonomous vehicles and may contribute to enhancing safety in both simulation and real-world applications. This research is intended solely for academic and simulation-based exploration. No deployable system is released. Any real-world deployment would require extensive additional testing, regulatory oversight, real-world validation, and continuous safety monitoring to ensure responsible use.